A 50 kg shell is fired from a cannon on Earth's surface with a potential energy of 196,000 J. What is the shell's maximum height?




1
Expert's answer
2021-02-04T11:43:40-0500

"PE=mgh\to h=\frac{PE}{mg}=\frac{196000}{50\cdot9.8}=400(m)" . Answer